OpenStudy (mathlegend):

the sum of 3 consecutive odd integers is 51. Find the numbers.

OpenStudy (mathlegend):

Using Algebra please

OpenStudy (anonymous):

i guess 11 11 + 13 + 15 = 39 nope too low try a higher number and i bet you will get it on the second try

OpenStudy (mathlegend):

I know the answer is 15, 17, 19

OpenStudy (anonymous):

x+x+1+x+2=51

OpenStudy (anonymous):

no need for any algebra, thinking is a human endeavour make an educated guess

OpenStudy (anonymous):

actually those are not consecutive odd integers try \(n+(n+2)+(n+4)=51\)

OpenStudy (mathlegend):

I did 3x+3=51 x = 16

OpenStudy (anonymous):

\(3n+6=51\) etc

OpenStudy (anonymous):

x=2n-1

OpenStudy (anonymous):

they are consecutive ODD integers, not consecutive integers that is why you need \(n,n+2,n+4\)

OpenStudy (mathlegend):

@satellite73 why did you do n+2 then n + 4?

OpenStudy (anonymous):

where n is odd

OpenStudy (mathlegend):

If I have n = 13 n + 2 = 15 n + 4 = 19 Don't I need 17?

OpenStudy (anonymous):

because n+1 and n+3 will be even if x+2 and x+4

OpenStudy (anonymous):

(2n + 1) + (2n + 3) + (2n + 5) = 51 is the equation you want because each term is guaranteed to be odd. And guessing is NOT the way to go. Extremely bad advice. That is NOT reflective of thinking like one of the helpers suggested. Being able to set up the proper equation shows and develops thinking. Bad advice, @satellite73 .

OpenStudy (anonymous):

\[(2n-1)+(2n+1)+(2n+3)=51\]\[3+6 n=51\]\[n=8 \]Plug in 8 for n to arrive at @tcarrollo10's solution.

OpenStudy (anonymous):

Here's the rest of my solution. I have this post and one other previous post. (2n + 1) + (2n + 3) + (2n + 5) = 51 -> 6n + 9 = 51 -> 6n = 42 -> n = 7 So, 2n + 1 = 15. The other 2 numbers are gotten the same way and are 17 and 19. This is the true way to get the answer. There is another "slick" way to get the answer (again NO GUESSING, that's just ridiculous). This "slick method" can be done in your head very quickly: take 51 and divide by 3. You get 17. Go up 2 on one of the 17's and down 2 on the other.